OpenStudy (anonymous):

need a check (4-z)z= Multiply and Simplify 4z^2?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

you have to multiply each term in the parentheses by \(z\)

OpenStudy (anonymous):

\(4-z\neq 4z\)

OpenStudy (anonymous):

4xz-zxz=4z

OpenStudy (anonymous):

or am I doing this wrong...

OpenStudy (anonymous):

\[(4-z)z=4\times z-z\times z\]

OpenStudy (anonymous):

no you are doing it right, but your answer is wrong

OpenStudy (anonymous):

\[4\times z=4z\] \[-z\times z=z^2\]

OpenStudy (anonymous):

you should get \[4z-z^2\]

OpenStudy (anonymous):

:( ok, i see now where i went wrong. i had it half right though lol